# Transit for All: Mapping Equitable Bike2Subway Connection using Region Representation Learning

### v1

@@ Abstract.
Ensuring equitable public transit access remains challenging, particularly in densely populated cities like New York City (NYC), where low-income and minority communities often face limited transit accessibility. Bike-sharing systems (BSS) can bridge these equity gaps by providing affordable first- and last-mile connections. However, strategically expanding BSS into underserved neighborhoods is difficult due to uncertain bike-sharing demand at newly planned (“cold-start”) station locations and limitations in traditional accessibility metrics that may overlook realistic bike usage potential. We introduce Transit for All (TFA), a spatial computing framework designed to guide the equitable expansion of BSS through three components: (1) spatially-informed bike-sharing demand prediction at cold-start stations using region representation learning that integrates multimodal geospatial data, (2) comprehensive transit accessibility assessment leveraging our novel weighted Public Transport Accessibility Level (wPTAL) by combining predicted bike-sharing demand with conventional transit accessibility metrics, and (3) strategic recommendations for new bike station placements that consider potential ridership and equity enhancement. Using NYC as a case study, we identify transit accessibility gaps that disproportionately impact low-income and minority communities in historically underserved neighborhoods. Our results show that strategically placing new stations guided by wPTAL notably reduces disparities in transit access related to economic and demographic factors. From our study, we demonstrate that TFA provides practical guidance for urban planners to promote equitable transit and enhance the quality of life in underserved urban communities.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
Strategic expansion of BSS into underserved communities requires careful planning for new station locations by considering (1) accurately predicting demand at newly planned (“cold-start”) stations and (2) effectively integrating predicted bike demand with existing transit accessibility measures. Traditional demand prediction methods, such as gravity-based models (Simini et al., 2021) and spatial interpolation (Liu and Oshan, 2022), rely on historical ridership data, limiting their usability for cold-start stations due to the absence of the ridership data. Recent deep-learning approaches (Liang et al., 2023) address this issue by integrating various types of built environment features. However, these methods use a simple concatenation of multimodal data, which may fail to capture the complex interactions among urban features, potentially affecting prediction accuracy. Moreover, existing accessibility metrics like Public Transport Accessibility Level (PTAL) (Transport for London, 2015; Adhvaryu and Kumar, 2021) primarily measure proximity and transit frequency but overlook realistic bike-to-transit usage potential, thus failing to provide practical insights into equitable accessibility.
@@
To overcome these limitations, we introduce Transit for All (TFA), a spatial computing framework designed to guide equitable BSS expansions. TFA has three components: (1) predicting bike demand at cold-start stations, (2) assessing realistic transit accessibility, and (3) recommending bike station placement location.
@@
First, to improve demand prediction at cold-start stations, TFA utilizes region representation learning (RRL), a multimodal geospatial learning approach that handles various geospatial data, including satellite imagery, points of interest (POIs), transit infrastructure, and neighborhood demographics. Unlike conventional methods that rely on the concatenation of raw features, RRL captures interactions among multiple urban datasets, addressing the complexity of urban dynamics between modalities. Second, to enhance traditional accessibility assessment methods, we introduce a novel weighted Public Transport Accessibility Level (wPTAL), which integrates predicted bike-sharing demand with the conventional PTAL metric. This integration captures realistic neighborhood-level differences in transit accessibility that traditional metrics may overlook. Third, based on our assessment via wPTAL, TFA strategically recommends new bike station locations that enhance equity by identifying underserved areas characterized by high predicted demand and low existing transit accessibility.
@@
We validate TFA’s practical effectiveness through a detailed case study in NYC. First, our analysis confirms that the RRL approach provides more accurate predictions for cold-start bike station demand compared to the baseline method that uses concatenated raw features. Second, our proposed wPTAL metric identifies significant yet previously unrecognized disparities in transit accessibility. Neighborhoods with predominantly Black, Hispanic, or medium-to-low-income populations indicate substantially lower wPTAL scores than major or high-income neighborhoods, meaning they have poorer accessibility compared to more affluent or centrally located neighborhoods. Finally, we quantitatively demonstrate that strategic station placement guided by TFA substantially reduce transit accessibility disparities across income and racial groups by using the Gini coefficient (Dorfman, 1979).
@@
  • C1 We introduce Transit for All, a spatial computing framework that predicts bike-sharing demand at cold-start stations without historical data by leveraging multimodal region representation learning.
  • C2: We propose a novel metric, the weighted Public Transport Accessibility Level (wPTAL), which integrates traditional transit accessibility metrics with predicted bike-sharing demand. This approach enables urban planners to strategically identify optimal locations for equitable bike station placements.
@@ 2.1. Distribution of Cold-start Stations in NYC
We define newly planned (“cold-start”) BSS stations, which have no previous records in a given month. Figure 1 shows the spatial distribution of Citi Bike stations across three periods in NYC. In 2017, 476 stations (shown in white in the 2017-2019 image in Figure 1) existed, predominantly in Manhattan and upper Brooklyn. Between 2017 and 2019, Citi Bike expanded, adding 341 new stations (shown in red), primarily extending into Queens and Brooklyn, increasing accessibility across residential neighborhoods.
@@ 3. Literature Review
Figure 3. Our spatial computing framework, Transit for All.
@@ 3.1. Mobility Infrastructure Planning Frameworks
Urban planners and policymakers increasingly use diverse transportation data to enhance equitable urban mobility. Existing frameworks commonly predict BSS demand using regression models informed by built-environment characteristics, evaluate pedestrian accessibility based on walking distances, and suggest new station locations (Chen et al., 2024; Beairsto et al., 2022; Mix et al., 2022). While these studies consider equity in station placement, they focus on pedestrian-focused accessibility, neglecting an integration with existing transit networks with BSS, such as subways or buses, limiting multimodal connectivity.
@@
In contrast to these studies, our Transit for All framework advances multimodal integration by (1) accurately predicting demand at cold-start bike stations using multimodal geospatial representation learning, (2) introducing a novel weighted transit accessibility metric (wPTAL) that combines predicted bike demand and existing public transit accessibility, and (3) strategically recommending equitable BSS expansions designed to improve first- and last-mile transit connectivity in underserved communities.
@@ 3.2. Transit Accessibility Measurements
Transit accessibility measurements quantify how easily residents can reach public transportation from given locations. Traditional metrics such as the Public Transport Accessibility Level (PTAL) evaluate accessibility based on proximity to transit infrastructure and scheduled transit frequencies (Transport for London, 2015).
Researchers adopt PTAL in cities beyond London (United Kingdom), including Lucknow (India) (Shah and Adhvaryu, 2016) and Hubli-Dharwad (India) (Adhvaryu and Kumar, 2021), supporting infrastructure planning in a variety of urban contexts. While PTAL identifies areas near transit stations, this approach does not incorporate actual or predicted travel demand. As a result, PTAL alone may not fully reflect the real-world transit needs of communities.
Recent approaches address this limitation by integrating observed transit demand into accessibility metrics. For instance, demand-weighted methods combine ridership data with infrastructure availability, helping to identify mismatches between transit supply and actual usage (Yen et al., 2023). However, these methods often neglect multimodal transportation connections, such as the connectivity between bike-sharing services and subway networks.
To bridge this gap, we introduce a weighted PTAL (wPTAL) metric that extends traditional PTAL by integrating factors such as bike travel time, waiting time, and transit frequency with predicted bike-sharing demand. By jointly considering infrastructure and expected usage, wPTAL enables planners to strategically place BSS stations and advance equitable transit accessibility across urban neighborhoods.
@@ 4. Transit for All
First, we apply region representation learning (RRL), utilizing multimodal data, such as satellite imagery, points-of-interest, demographics, and existing transit infrastructure, to generate region vector and predict bike demand with a supervised predictor. Second, we introduce a novel accessible metic, the weighted Public Transport Accessibility Level (wPTAL), which integrates predicted ridership with transit accessibility to better evaluate station locations. Finally, we recommend future bike station placement locations based on wPTAL scores, enabling planners to address transit equity gaps in underserved communities.
@@ 4.1.1. Region Representation Learning
RRL encodes geospatial data into compact embeddings representing geographic areas. We use the Cross-modal Knowledge Injected Embedding (CooKIE[Uncaptioned image](Namgung et al., 2025) to integrate multimodal data, such as satellite imagery, urban infrastructure, and socio-demographics into unified region embeddings. Existing RRL approaches often concatenate raw features directly, which may overlook the complementary relationships between modalities. For example, satellite imagery captures spatial patterns, socio-demographics describe community characteristics, and infrastructure reflects land functionality. Direct concatenation may miss these complementary multimodal relationships.
@@
where $x_{i}^{1},x_{i}^{2},x_{i}^{3}$ denote multimodal input features from image data, urban infrastructure, and socio-demographics (see details in Section 5.2), and $x_{i}^{c}$ denotes the final learned region embedding used for predicting bike demand at station $i$.
@@ 4.1.2. Predictor
After extracting the learned region representation via CooKIE, we feed it into our predictor, Spatial-MGAT (Liang et al., 2023). Bike demand at a newly planned station is often influenced by two types of existing stations: (1) nearby stations with similar travel patterns, and (2) stations that share similar built-environment characteristics.
To capture these spatial relationships, Spatial-MGAT constructs two localized graphs for each cold-start station $i$:
  • $G^{p}_{i,m}$: connects station $i$ to the $k$ geographically nearest stations.
  • $G^{b}_{i,m}$: connects station $i$ to the $k$ most similar stations based on built-environment features (e.g., points-of-interest, transit infrastructure, and socio-economic attributes).
From these graphs, we derive spatial-interaction vectors $s^{(p)}_{i,m}$ and $s^{(b)}_{i,m}$ by aggregating neighbor embeddings via learned attention:
@@
Here, the neighbor sets $\mathcal{N}^{(p)}_{i,m}$ and $\mathcal{N}^{(b)}_{i,m}$ correspond to the geographical proximity graph $G^{p}_{i,m}$ and the built-environment similarity graph $G^{b}_{i,m}$.
The neighbor embeddings $h_{j,m}$ are weighted by attention coefficients $\epsilon^{(*)}_{ij,m}$, and $h_{j,m}$ is derived by $\text{MLP}([x_{j}^{c}\parallel x_{j,m}^{u}])$, where the concatenation ($\parallel$) of region embedding from CooKIE with monthly urban context features $x_{j,m}^{u}$.
We calculate the attention coefficients $\epsilon^{(*)}_{ij,m}$ for each localized graph $*\in\{p,b\}$ through a softmax over pairwise attention scores:
@@
Finally, we concatenate the station’s CooKIE’s embedding $x_{i}^{c}$, monthly urban context $x_{i,m}^{u}$, temporal features $t_{i,m}$, and spatial interaction vectors from the two graphs to predict the bike demand with a two-layer feed-forward neural network:
@@ 4.2. Assess Public Transit Accessibility
To better reflect realistic transit accessibility and guide equitable infrastructure planning, we propose the Weighted Public Transport Accessibility Level (wPTAL) by integrating predicted bike demand from Spatial-MGAT. Traditional metrics like the Public Transport Accessibility Level (PTAL), developed by Transport for London (Transport for London, 2015), primarily measure accessibility based on transit proximity and service frequency.
Public Transport Accessibility Level (PTAL): PTAL quantifies accessibility by aggregating subway access frequency via the Equivalent Doorstep Frequency (EDF), which integrates bike travel time and subway waiting periods. EDF is calculated by:
@@
(9) $PTAL_{i}=\max_{j}(EDF_{ij})+0.5\times\sum_{j\neq j_{max}}EDF_{ij}.$
However, PTAL alone does not account for potential bike demand usage at cold-start stations, which is obtained from Section 4.1.
Weighted Public Transport Accessibility Level (wPTAL): Our proposed metric enhances traditional PTAL by incorporating predicted bike demand ($\hat{y}_{i,m}$) from Spatial-MGAT. Traditional PTAL solely considers accessibility based on proximity and service frequency without considering the potential utilization by local residents. For example, consider two neighborhoods with identical PTAL scores. The neighborhood with higher predicted bike usage would likely experience greater practical benefits from the bike-sharing station compared to the neighborhood with lower predicted usage, as a larger number of residents are expected to use bikes to connect to transit services. Thus, wPTAL captures not only how accessible a station is theoretically (transit proximity and frequency) but also how beneficial or useful it will realistically be to the community (predicted demand):
(10) $wPTAL_{i}=PTAL_{i}\times\hat{y}_{i,m},$
where $\hat{y}_{i,m}$ is predicted bike demand at cold-start station $i$, and $PTAL_{i}$ is the PTAL score at $i$.

### v2

@@ Abstract.
Ensuring equitable public transit access remains challenging, particularly in densely populated cities like New York City (NYC), where low-income and minority communities often face limited transit accessibility. Bike-sharing systems (BSS) can bridge these equity gaps by providing affordable first- and last-mile connections. However, strategically expanding BSS into underserved neighborhoods is difficult due to uncertain bike-sharing demand at newly planned (“cold-start”) station locations and limitations in traditional accessibility metrics that may overlook realistic bike usage potential. We introduce Transit for All (TFA), a spatial computing framework designed to guide the equitable expansion of BSS through three components: (1) spatially-informed bike-sharing demand prediction at cold-start stations using region representation learning that integrates multimodal geospatial data, (2) comprehensive transit accessibility assessment leveraging our novel weighted Accessibility Level (wAL) by combining predicted bike-sharing demand with conventional transit accessibility metrics, and (3) strategic recommendations for new bike station placements that consider potential ridership and equity enhancement. Using NYC as a case study, we identify transit accessibility gaps that disproportionately impact low-income and minority communities in historically underserved neighborhoods. Our results show that strategically placing new stations guided by wAL notably reduces disparities in transit access related to economic and demographic factors. From our study, we demonstrate that TFA provides practical guidance for urban planners to promote equitable transit and enhance the quality of life in underserved urban communities.
@@
journalyear: 2025copyright: acmlicensedconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances in Geographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3748636.3764176isbn: 979-8-4007-2086-4/2025/11ccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
Strategic expansion of BSS into underserved communities requires careful planning for new station locations by considering (1) accurately predicting demand at newly planned (“cold-start”) stations and (2) effectively integrating predicted bike demand with existing transit accessibility measures. Traditional demand prediction methods, such as gravity-based models (Simini et al., 2021) and spatial interpolation (Liu and Oshan, 2022), rely on historical ridership data, limiting their usability for cold-start stations due to the absence of the ridership data. Recent deep-learning approaches (Liang et al., 2023) address this issue by integrating various types of built environment features. However, these methods use a simple concatenation of multimodal data, which may fail to capture the complex interactions among urban features, potentially affecting prediction accuracy. Moreover, existing accessibility metrics like Public Transit Accessibility Level (PTAL) (Transport for London, 2015; Adhvaryu and Kumar, 2021) primarily measure proximity and transit frequency but overlook realistic bike-to-transit potential usage, thus failing to provide practical insights into equitable accessibility.
@@
To overcome these limitations, we introduce Transit for All (TFA), a spatial computing framework designed to guide equitable BSS expansions. TFA has three components: (1) predicting bike demand at cold-start stations, (2) assessing realistic transit accessibility, and (3) recommending bike station placement.
@@
First, to improve demand prediction at cold-start stations, TFA utilizes multimodal region representation learning (RRL), that handles various geospatial data, including satellite imagery, point of interest (POIs), transit infrastructure, and neighborhood demographics. Unlike conventional methods that rely on the concatenation of raw features, RRL captures interactions among multiple datasets, addressing the complexity of urban dynamics between modalities. Second, to enhance traditional accessibility assessment methods, we introduce a novel weighted Accessibility Level (wAL), which integrates predicted bike-sharing demand with the conventional PTAL metric. This integration captures realistic neighborhood-level differences in transit accessibility that traditional metrics may overlook. Third, based on our assessment via wAL, TFA strategically recommends new bike station locations that enhance equity by considering underserved areas characterized by high predicted demand and low existing transit accessibility.
@@
We validate TFA’s effectiveness through a detailed case study in NYC. First, our analysis confirms that the RRL approach provides more accurate predictions for cold-start bike station demand compared to the baseline method that uses concatenated raw features. Second, our proposed wAL metric identifies significant yet previously unrecognized disparities in transit accessibility. Neighborhoods with predominantly Black, Hispanic, or medium-to-low-income populations indicate substantially lower wAL scores than major or high-income neighborhoods, meaning they have poorer accessibility compared to more affluent or centrally located neighborhoods. Finally, we quantitatively demonstrate that strategic station placement guided by TFA substantially reduce transit accessibility disparities across income and racial groups by using the Gini coefficient (Dorfman, 1979).
@@
  • C1 We introduce Transit for All, a spatial computing framework that predicts bike-sharing demand at cold-start stations by leveraging multimodal region representation learning.
  • C2: We propose a novel metric, the weighted Accessibility Level (wAL), which integrates traditional transit accessibility metrics with predicted bike-sharing demand. This approach enables urban planners to strategically identify optimal locations for equitable bike station placements.
@@ 2.1. Distribution of Cold-start Stations in NYC
We define newly planned (“cold-start”) BSS stations, which have no previous records in a given month. Figure 1 shows the spatial distribution of Citi Bike stations across three periods in NYC. In 2017, 476 stations (shown in white in the 2017-2019 image in Figure 1) existed, predominantly in Manhattan and upper Brooklyn. Between 2017 and 2019, Citi Bike expanded, adding 341 new stations (shown in red), primarily expanding to Queens and Brooklyn, increasing accessibility across residential neighborhoods.
@@ 3. Literature Review
Figure 3. An overview of our Transit for All (TFA) framework. TFA consists of four components: (1) Learning region representaion via CooKIE, (2) Predicting bike demand at each cold-start station using Spatial-MGAT, (3) Assessing bike to transit accessibility by computing wAL, and (4) Recommending future bike station placement.
@@ 3.1. Mobility Infrastructure Planning Frameworks
Urban planners and policymakers increasingly use diverse transportation data to enhance equitable urban mobility. Existing frameworks commonly predict BSS demand using regression models informed by built environment characteristics, evaluate pedestrian accessibility based on walking distances, and suggest new station locations (Chen et al., 2024; Beairsto et al., 2022; Mix et al., 2022). While these studies consider equity in station placement, they focus on pedestrian-focused accessibility, neglecting an integration with existing transit networks with BSS, such as subways or buses, limiting multimodal connectivity.
@@
In contrast to these studies, our Transit for All framework advances multimodal integration by (1) predicting demand at cold-start bike stations using multimodal representation learning, (2) introducing a novel weighted transit accessibility metric (wAL) that combines predicted bike demand and existing public transit accessibility, and (3) strategically recommending equitable BSS expansions designed to improve first- and last-mile transit connectivity in underserved communities.
@@ 3.2. Transit Accessibility Measurements
Transit accessibility measurements quantify how easily residents can reach public transportation from given locations. Traditional metrics such as the Accessibility Level (AL) evaluate accessibility based on proximity to transit infrastructure and scheduled transit frequencies (Transport for London, 2015).
Researchers adopt AL in cities beyond London (United Kingdom), including Lucknow (India) (Shah and Adhvaryu, 2016) and Hubli-Dharwad (India) (Adhvaryu and Kumar, 2021), supporting infrastructure planning in a variety of urban contexts. While AL identifies areas near transit stations, this approach does not incorporate actual or predicted travel demand. As a result, AL alone may not fully reflect the real-world transit needs of communities.
Recent approaches address this limitation by integrating observed transit demand into accessibility metrics. For instance, demand weighted methods combine ridership data with infrastructure availability, helping to identify mismatches between transit supply and actual usage (Yen et al., 2023). However, these methods often neglect multimodal transportation connections, such as the connectivity between bike-sharing services and subway networks.
To bridge this gap, we introduce a weighted AL (wAL) metric that extends traditional AL by integrating factors such as bike travel time, waiting time, and transit frequency with predicted bike-sharing demand. By jointly considering infrastructure and expected usage, wAL enables planners to strategically place BSS stations and advance equitable transit accessibility across urban neighborhoods.
@@ 4. Transit for All
First, we apply region representation learning (RRL), utilizing multimodal data, such as satellite imagery, points-of-interest, demographics, and existing transit infrastructure, to generate region vector and predict bike demand with a supervised predictor. Second, we introduce a novel accessible metric, the weighted Accessibility Level (wAL), which integrates predicted ridership with transit accessibility to better evaluate station locations. Finally, we recommend future bike station placement locations based on wAL scores, enabling planners to address transit equity gaps in underserved communities.
@@ 4.1.1. Region Representation Learning
RRL encodes geospatial data into compact embeddings representing geographic areas. We use the Cross-modal Knowledge Injected Embedding (CooKIE[Uncaptioned image](Namgung et al., 2025) to integrate multimodal data, such as satellite imagery, urban infrastructure, and socio-demographics into unified region embeddings. Existing RRL approaches often concatenate raw features directly, which may overlook the complementary relationships between modalities. For example, satellite imagery captures spatial patterns, socio-demographics describe community characteristics, and infrastructure reflects land functions. Direct concatenation may miss these complementary multimodal relationships.
@@
where $x_{i}^{1},x_{i}^{2},x_{i}^{3}$ denote multimodal input features from satellite imagery, urban infrastructure, and socio-demographics (see details in Section 5.2), and $x_{i}^{c}$ denotes the final learned region embedding used for predicting bike demand at station $i$.
@@ 4.1.2. Predictor
After extracting the learned region representation via CooKIE, we feed it into our predictor, Spatial-MGAT (Liang et al., 2023). Bike demand at a newly planned station is often influenced by two types of existing stations: (1) nearby stations with similar travel patterns, and (2) stations that share similar built environment characteristics.
To capture these dependencies, we denote by $m$ the month (time step) and construct two localized graphs for each target station $i$:
  • $G^{p}_{i,m}$: a proximity-based graph that links the target station $i$ to its $k$ geographically nearest stations ($p$ denotes geographical proximity).
  • $G^{b}_{i,m}$: a built environment similarity graph that connects stations $i$ to its $k$ most similar stations based on built-environment characteristics (e.g., POIs, transit infrastructure, and socio-economic attributes), where $b$ denotes built environment similarity.
For each graph $G^{*}_{i,m}(*\in\{p,b\})$, we denote by $\mathcal{N}^{(*)}_{i,m}$ the set of neighbor stations directly connected to the target stations $i$. From these graphs, we derive spatial-interaction vectors $s^{(p)}_{i,m}$ and $s^{(b)}_{i,m}$ by aggregating neighbor embeddings via learned attention:
@@
Here,
  • $h_{j,m}$ is the embedding of a neighbor station $j$ in month $m$, computed as
    (6) $h_{j,m}=\text{MLP}([x_{j}^{c}\parallel x_{j,m}^{u}]),$
    where $x_{j}^{c}\parallel x_{j,m}^{u}]$ concatenates the learned region embedding $x_{j}^{c}$ from CooKIE with the monthly urban context features $x_{j,m}^{u}$.
  • $\epsilon^{(*)}_{ij,m}$ is the attention coefficient that measures the relative importance of neighbor $j$. Specifically,
@@
Finally, we concatenate each station’s CooKIE’s embedding $x_{i}^{c}$, monthly urban context $x_{i,m}^{u}$, temporal features $t_{i,m}$, and spatial interaction vectors from the two graphs to predict the bike demand with a two-layer feed-forward neural network:
@@ 4.2. Assess Public Transit Accessibility
To better reflect realistic transit accessibility and guide equitable infrastructure planning, we propose the weighted Accessibility Level (wAL) by integrating predicted bike demand from Spatial-MGAT. Traditional metrics like the Accessibility Level (AL), developed by Transport for London (Transport for London, 2015), primarily measure accessibility based on transit proximity and service frequency.
Accessibility Level (AL): AL quantifies accessibility by aggregating subway access frequency via the Equivalent Doorstep Frequency (EDF), which integrates bike travel time and subway waiting periods. EDF is calculated by:
@@
(10) $AL_{i}=\max_{j}(EDF_{ij})+0.5\times\sum_{j\neq j_{max}}EDF_{ij}.$
However, AL alone does not account for potential bike demand at cold-start stations, which is obtained from Section 4.1. For example, two stations may have the same AL, but if one has much lower bike demand, its accessibility will translate into far less actual use.
weighted Accessibility Level (wAL): Our proposed metric enhances traditional AL by incorporating predicted bike demand ($\hat{y}_{i,m}$) from Spatial-MGAT. Traditional AL solely considers accessibility based on proximity and service frequency without considering the potential utilization by local residents. For example, consider two neighborhoods with identical AL scores. The neighborhood with higher predicted bike usage would likely experience greater practical benefits from the bike-sharing station compared to the neighborhood with lower predicted usage, as a larger number of residents are expected to use bikes to connect to transit services. Thus, wAL captures not only how accessible a station is theoretically (transit proximity and frequency) but also how beneficial or useful it will realistically be to the community (predicted demand):
(11) $wAL_{i}=AL_{i}\times\hat{y}_{i,m},$
where $\hat{y}_{i,m}$ is predicted bike demand at cold-start station $i$, and $AL_{i}$ is the AL score at $i$.